Exercise 4-1 (Video) Saddle Inc. has two types of handbags: standard and custom. The costing and wants to see how the results would differ if this companys operations controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the Standard Custom Direct labor costs Machine hours Setup hours $50,000 100,000 1,200 420 1,200 90 Total estimated overhead costs are $299,100. Overhead cost allocated to the machining activity cost pool is $192,000, and $107,100 is allocated to the machine setup activity cost pool Compute the overhead rate using the traditional (plantwide) approach. (Round answer to 2 decimal places, e9. 12.25.) Predetermined overhead rate % of direct labor cost Compute the overhead rates using the activity-based costing approach. Machining per machine hour Machine setup per setup hour
Compute the overhead rates using the activity-based costing approach Machining Machine setup per machine hour per setup hour LINK TO TEXT LINK TO TEXT VIDEO SIMILAR EXERCISE Determine the difference in allocation between the two approaches Traditional costing Standard Custom Activity-based costing Standard Custom Click If you would like to Show Work for this question: Open Show.Work LINK TO TEXYLN
Answer #1

(i)

Total estimated overhead costs = $299,100

Total direct labor costs = $150,000

Predetermined overhead rate = Total estimated overhead costs/Total direct labor costs

= 299,100/150,000

= 199.4%

(ii)

Overhead rates using Activity based costing approach

Machining = Overhead cost allocated to machining activity cost pool/Number of machine hours

= 192,000/2,400

= $80 per machine hour

Machine set up = Overhead cost allocated to set up activity cost pool/Number of set up hours

= 107,100/510

= $210 per set up hour

(iii)

Difference in allocation between two approaches

Traditional costing

Standard = 50,000 x 199.4%

= $99,700

Custom = 100,000 x 199.4%

= $199,400

Activity based costing

Standard = 1,200 x 80 + 90 x 210

= 96,000 + 18,900

= $114,900

Custom = 1,200 x 80 + 420 x 210

= 96,000 + 88,200

= $184,200

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.